Suppose that an automobile manufacturer advertises that its new hybrid car has a mean gas mileage of 50 miles per gallon. You take a simple random sample of n = 30 hybrid vehicles and test their gas mileage. You find that in this sample, the average is ̄x = 47 miles per gallon with a standard deviation of s = 5.5 miles per gallon. Does this indicate that the advertiser’s statement is true or false? Let a=.0.05
1
Expert's answer
2019-12-03T10:55:51-0500

We have "\mu_0=50"

For the sample, we have

"n=30, \bar{X}=47" and

 "S=5.5"

Null Hypothesis, "H_0:\mu=\mu_0"

Alternate hypothesis, "H_1:\mu\not=\mu_0"


Test statistics, "t=(\bar{x}-\mu_0)/(S/\sqrt{n})"

"=(47-50)/(5.5/\sqrt{30})"

"=-3\/1.004=-2.9875"

Since the population standard deviation is unknown and the sample size is also small, we will use t-test.

Degree of freedom (DOF) for this t-test will be equal to "n-1=30-1=29" .

The t-critical values for a two-tailed test, for a significance level of "\alpha= 0.05" and "DOF=29" will be

"t_c<-2.045" and "t_c>2.045".

As we can see that our test statistics lies in the critical region, we will have to reject the null hypothesis.

So, we can say that at "\alpha =0.05" the advertiser's statement is false.